OpenStudy (anonymous):

what is the vertex of the graph of y=-4(x-1)(x+5)

sam (.sam.):

expand it first

OpenStudy (anonymous):

uuuuuuuum (-4x_4)(x+5)

OpenStudy (anonymous):

?

OpenStudy (callisto):

I assume _ is + and then expand (-4x+4)(x+5)

OpenStudy (anonymous):

no its - sorry

OpenStudy (callisto):

Oh.. then you were wrong..

OpenStudy (callisto):

y=-4(x-1)(x+5) = [(-4)x -(-4)(1)] (x+5) = (-4x + 4)(x+5)

OpenStudy (callisto):

Got it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but what do i do after that

OpenStudy (callisto):

Now, expand (-4x+4)(x+5)

OpenStudy (anonymous):

so combine? them

OpenStudy (callisto):

Expand them...

OpenStudy (callisto):

(-4x+4)(x+5) = -4x(x+5) + 4(x+5) = ?

OpenStudy (anonymous):

-4x^2-20x+4x+5 -4x^2-16x+5

OpenStudy (callisto):

Nope, the constant term is incorrect.

OpenStudy (anonymous):

ummm

OpenStudy (callisto):

(-4x+4)(x+5) = -4x(x+5) + 4(x+5) = -4x(x) -4x(5) + 4x + 4(5) =?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

and 5

OpenStudy (callisto):

._. Nope.... Show your workings please

OpenStudy (callisto):

._. Perhaps I show it to you... -4x(x+5) + 4(x+5) = -4x^2 -20x + 4x +20 = -4x^2 -16x + 20 You need to know what it is 4 times 5, which is equal to 20, for the constant term.. Got it?

OpenStudy (anonymous):

ok so the constant term is 20?

OpenStudy (callisto):

Yes... Now, what we need to do is to completing square.. y = -4x^2 -16x + 20 = -4(x^2 +4x) + 20 = -4(x^2 + 4x +4 -4) +20 = -4 (x+2)^2 + (-4)(-4) + 20 = ??

OpenStudy (anonymous):

YOu don't need to expand it to find the vertex!! since this quadratic is in "intercept form", identify the 2 x-intercepts! They are 1 and -5... the axis of symmetry is the midpoint of these 2 numbers: -2..... Now, plug -2 into the original equation to determine the vertex!! (-2, 36)

OpenStudy (anonymous):

wooooooooooah

OpenStudy (anonymous):

There are other ways to find the vertex, but this is the easiest way (for this instance)

OpenStudy (callisto):

:| Haven't thought of that... I'm stupid :(